# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ B.1 Datasets
Table 6: Dataset Statistics.
@@
Size of Anchor Set ${\mathcal{D}}$ 640 (100%) 559 (50%) 562 (25%) 564 (25%) 743 (15%) 942 (10%) 2,065 (10%)
@@
B.2 Evaluation Metrics
Negative Log-Likelihood (NLL) and Expected Calibration Error (ECE (Naeini et al., 2015)) are key metrics for uncertainty estimation. For a model $P_{\bm{\theta}}$ and test dataset $\{{\bm{x}}_{n},y_{n}\}_{n=1}^{N}$, NLL penalizes models that assign low probabilities to correct labels, and is defined as:
@@ B.2 Evaluation Metrics
where $\text{acc}(B_{m})=\nicefrac{{1}}{{|B_{m}|}}\sum_{i\in B_{m}}\mathbbm{1}{(\widehat{y}_{i}=y_{i})}$ is the average accuracy and $\text{conf}(B_{m})=\nicefrac{{1}}{{|B_{m}|}}\sum_{i\in B_{m}}P(\widehat{y}_{i})$ is the average confidence in bin $B_{m}$. We use bin size $|B_{m}|=15$ throughout our experiments.
B.3 Optimal $\sigma_{q}^{*}$ of TFB Variants
TFB converts the end-to-end training using variational inference to a variance maximization problem. It is done through trimming down the set of trainable parameters to a single variable $\sigma_{q}$. How to use this single variable $\sigma_{q}$ to define a better distribution family remains an intriguing problem for future research.
In Sec. 5.3, to show more possibilities of such, we extend TFB to other families of Gaussians for modeling the variational distributions of ${\bm{W}}$. Specifically, we consider: (i) Full-Rank Isotropic Gaussian (FR, ${\bm{\Sigma}}_{q}=\sigma_{q}^{2}{\bm{I}}$), and (ii) Constant Low-Rank Standard Deviation (C-STD, ${\bm{\Omega}}=[\Omega_{ij}=\sigma_{q}]$).
For fair comparison, we adopt the same optimal $\sigma_{q}^{*}$ search protocol of TFB as decribed in Sec. 5.1. In Table 7, we show the optimal setting of $\sigma_{q}^{*}$ with adaptive performance change tolerance $\epsilon=1\%$ or $\epsilon=0.5\%$ searched for all three variational distribution families. Additionally, we show the upper-bound performance of each of them by directly tuning $\sigma_{q}$ on the test set, aiming to minimize the ECE.
Table 7: Searched $\sigma_{q}^{*}$ of Training-Free Bayesianization methods, where “*” represents that $\sigma_{q}^{*}$ is tuned for optimal calibration (ECE).
     Base Model           $\sigma_{q}^{*}$      WG-S      ARC-C      ARC-E      WG-M      OBQA      BoolQ
     MLE      TFB* (C-STD)      0.13      0.10      0.12      0.10      0.10      0.10
     TFB* (FR)      0.003      0.0025      0.003      0.003      0.003      0.003
     TFB* (Ours)      0.04      0.024      0.018      0.02      0.012      0.008
     TFB (Ours)      0.0216      0.0201      0.0211      0.0156      0.0145      0.0112
     MAP      TFB* (C-STD)      0.12      0.11      0.138      0.11      0.1      0.1
     TFB* (FR)      0.003      0.0025      0.003      0.0025      0.003      0.003
     TFB* (Ours)      0.03      0.025      0.016      0.022      0.013      0.009
     TFB (Ours)      0.0277      0.0222      0.0226      0.0202      0.0151      0.0150
     BLoB-Mean      TFB* (C-STD)      0.1      0.18      0.18      0.1      0.1      0.09
     TFB (C-STD)      0.1473      0.1817      0.1518      0.1261      0.0885      0.1422
     TFB* (FR)      0.004      0.004      0.003      0.003      0.003      0.003
     TFB (FR)      0.0037      0.0006      0.0034      0.0034      0.0035      0.0034
     TFB* (Ours)      0.03      0.016      0.01      0.015      0.006      0.007
     TFB (Ours)      0.0273      0.0137      0.0163      0.0194      0.0064      0.0157
B.4 Training
@@
We report the mean and standard deviation of all experimental results calculated over three random seeds. For all training processes in our experiments, we employ the AdamW optimizer. The learning rate follows a linear decay schedule with a warmup ratio of 0.06 and a maximum value of $41e-$ 4. The batch size is set to 4, and the maximum sentence length is limited to 300 tokens. Additionally, the LoRA configuration includes LoRA $\alpha=16$ and LoRA $r=8$. PiSSA (Meng et al., 2024) follows the exact same configuration as the LoRA’s. For VeRA (Kopiczko et al., 2023), due to its characteristic of shared weights across different layers, we set its rank $r=256$ and learning rate to $35e-$ 3 for the MLE training on the combined dataset.
Appendix C Additional Experimental Results
C.1 TFB Variants: $\epsilon$s and Optimal Calibration
In Table 1 (Sec. 5.2), we present TFB results using adaptive $\epsilon$ (1% or 0.5%) based on LoRA’s overfitting status. Table 8 shows comprehensive results for both $\epsilon$ values, along with an Optimal ECE where $\sigma_{q}$ is optimized on test data for best ECE. This bound demonstrates TFB’s potential when a calibration dataset is available, though this reduces it to a degraded calibration method.

### v2

@@ E.1 Datasets
Table 6: Dataset Statistics. The size of the Anchor Set ${\mathcal{D}}$ is used in Table 1, 2 and 14.
@@
Size of Anchor Set ${\mathcal{D}}$ 500 (78%) 500 (45%) 500 (22%) 500 (22%) 500 (10%) 500 (5%) 500 (2%)
@@
E.2 Evaluation Metrics
Negative Log-Likelihood (NLL) and Expected Calibration Error (ECE [45]) are key metrics for uncertainty estimation. For a model $P_{\bm{\theta}}$ and test dataset $\{{\bm{x}}_{n},y_{n}\}_{n=1}^{N}$, NLL penalizes models that assign low probabilities to correct labels, and is defined as:
@@ E.2 Evaluation Metrics
where $\text{acc}(B_{m})=\nicefrac{{1}}{{|B_{m}|}}\sum_{i\in B_{m}}\mathbbm{1}{(\widehat{y}_{i}=y_{i})}$ is the average accuracy and $\text{conf}(B_{m})=\nicefrac{{1}}{{|B_{m}|}}\sum_{i\in B_{m}}P(\widehat{y}_{i})$ is the average confidence in bin $B_{m}$. We use bin size $|B_{m}|=15$ throughout this paper.
E.3 Searched $\sigma_{q}$ of TFB
We report the searched $\sigma_{q}^{*}$ using Algorithm A in Table 7, where the reported values are the mean values of three random seeds.
Table 7: Searched $\sigma_{q}^{*}$ of TFB using Algorithm A.
     Base Model      WG-S      ARC-C      ARC-E      WG-M      OBQA      BoolQ
     MLE      0.004500      0.003917      0.004500      0.004354      0.003771      0.004063
     MAP      0.004500      0.003479      0.003188      0.004208      0.003917      0.005083
     BLoB-Mean      0.005813      0.005229      0.005229      0.006250      0.006250      0.005958
E.4 Bayesianization (Training)
@@
Shared Configuration. We report the mean and standard deviation of all experimental results calculated over three random seeds. For all training processes in our experiments, we employ the AdamW optimizer. The learning rate follows a linear decay schedule with a warmup ratio of 0.06 and a maximum value of $42e-$ 4. The batch size is set to 4, and the maximum sentence length is limited to 300 tokens. The LoRA configuration includes LoRA $\alpha=16$ and LoRA $r=8$. PiSSA [41] follows the exact same configuration as the LoRA’s. For VeRA [35], due to its characteristic of shared weights across different layers which enables higher-rank setting with the same memory efficiency, we set its rank to $r=256$ and learning rate to $35e-$ 3 for the MLE training on the combined dataset.
Baseline Configuration. The baseline configuration mainly follows BLoB [61]. MLE follows the standard LoRA implementation. For MAP, we implement it with a weight decay rate of $51e-$ 5. MCD consists of an ensemble of 10 LoRAs with a dropout rate of $p=0.1$. For ENS, we fine-tune 3 LoRAs independently and combine them by averaging their logits during evaluation. We implement LAP and apply it to the MAP checkpoints. For BBB and BLoB, we use the default settings from Bayesian-Torch library [36], applying Bayesianization only to the ${\bm{A}}$ matrix. During training, the number of samples is set to $K=1$ for both BBB and BLoB. At test time, we use $N=10$ samples, matching the configuration of TFB.
TFB Configuration. We randomly sample unlabeled training data points to construct the anchor dataset ${\mathcal{D}}=\{{\bm{x}}_{i},\widehat{y}_{i}\}_{i\in[M]}$ where $\widehat{y}_{i}$ is the pseudo-label generated by the given LoRA adapter before Bayesianization; the anchor dataset size $M=500$ is fixed for all the datasets. We use NLL as the metric $l$ and set the performance change tolerance $\epsilon$ to 0.3% of relative performance change for all the datasets. To determine the optimal $\sigma_{q}^{*}$, we perform a 5-step binary search with the initial range of $[0.001,0.015]$ using Algorithm A. Similar to the other baseline methods, the final results of TFB are reported as averages across three random seeds using $\sigma_{q}^{*}$.
Appendix F Additional Experimental Results
We present additional experimental results in this section. Due to space constraints (and large table size), we defer several detailed tables to the end of this section rather than presenting them alongside the corresponding analyses.
Figure 2: Uncertainty estimation of TFB on a toy regression task. The true cubic function $y=x^{3}$ (GREEN) and noisy training samples (GRAY) are shown alongside predictions from a deterministic MLP baseline (RED) and our TFB (BLUE). The blue shaded region represents ±1 standard deviation from TFB predictions with $\sigma_{q}=1.0$. TFB effectively captures predictive uncertainty, showing low variance in data-dense regions ($x\in[0,2]$) and increasing uncertainty for inputs outside the training distribution ($x\notin[-4,4]$), with uncertainty proportional to distance from the training domain.
F.1 TFB as a General Bayesian Neural Network: A Visual Study
We demonstrate that TFB functions as a general Bayesian Neural Network (beyond its application to the LoRA adapters of LLMs) by evaluating its uncertainty estimation capabilities on a toy regression dataset, as illustrated in Fig. 2. We follow the regression task framework [21, 56] with the following specifications:
  • Input Features: $\{{\bm{x}}_{i}\sim U[-4,4]\}_{i\in[20]}$ uniformly sampled from the interval $[-4,4]$.
  • True Function: $y=x^{3}$ (plotted in GREEN).
  • Noisy Labels: ${\bm{y}}_{i}={\bm{x}}_{i}^{3}+9\cdot\epsilon_{i}$ where $\epsilon_{i}\sim{\mathcal{N}}(0,1)$, representing the true function with added Gaussian noise. These data points appear as GRAY crosses.
  • MLE Baseline: A two-layer MLP with 16 hidden neurons fit to the sampled dataset, providing a deterministic baseline without uncertainty quantification. The MLE predictions are shown in RED.
  • MLE Training Configuration: Adam optimizer [31] with learning rate 0.1, trained for 1000 steps until convergence.
  • TFB Implementation: We Bayesianize the MLE baseline using TFB with full-rank noise, defining the variational distribution as $q({\bm{W}}|{\bm{W}}_{0},\sigma_{q})=\prod_{ij}{\mathcal{N}}(W_{ij}|W_{0,ij},\sigma_{q}^{2})$. For inference, we draw $N=10$ samples from the variational weight distribution and plot the average prediction as a BLUE curve. The shaded region represents ±$\pm 1$ standard deviation of the sampled predictions. We evaluate multiple settings of $\sigma_{q}\in[0.1,0.3,0.6,1.0,1.5]$ and report the average squared predictive difference between TFB and MLE in the figure.
As Fig. 2 demonstrates, with an appropriately calibrated standard deviation of the approximate posterior ($\sigma_{q}=1.0$ in this case), TFB effectively captures predictive uncertainty: in regions with dense data sampling (${\bm{x}}\in[0,2]$), TFB produces low predictive uncertainty; conversely, for inputs outside the training domain (${\bm{x}}\notin[-4,4]$), the predictive uncertainty increases proportionally with distance from the training distribution.
F.2 Computational Complexity Analysis
We compare the computational efficiency of TFB and BLoB during the process of Bayesianization in Table 8. We also report the computational cost of the standard LoRA fine-tuning as reference. All three methods are evaluated on the configurations detailed in Appendix E.4. For LoRA and BLoB, the evaluation of running time and maximum GPU memory is based on fine-tuning for 5 epochs. TFB uses a fixed number of 500 training examples to search for $\sigma_{q}^{*}$ across all datasets, and performs binary search for at most 5 rounds (sequentially).
Table 8: A comparison of running time and maximum GPU memory cost between TFB and BLoB during the process of Bayesianizatioin. The experiments are conducted on the single NVIDIA A100 GPU setting. The subscripts in the table calculate the relative cost of a method compared to that of LoRA, a non-Bayesian baseline method. RED and GREEN represent worse and better efficiency, respectivley. Note that varying batch sizes do not impact the performance of TFB, as Algorithm A is independent of gradient and batch size
Method Batch Size Datasets
WG-S ARC-C ARC-E WG-M OBQA BoolQ
Time (s) Mem. (MB) Time (s) Mem. (MB) Time (s) Mem. (MB) Time (s) Mem. (MB) Time (s) Mem. (MB) Time (s) Mem. (MB)
LoRA 4 338 (0.00x) 12,894 (0.00x) 632 (0.00x) 19,762 (0.00x) 1,238 (0.00x) 18,640 (0.00x) 1,339 (0.00x) 13,164 (0.00x) 2,692 (0.00x) 17,208 (0.00x) 6,489 (0.00x) 29,450 (0.00x)
BLoB 4 371 (1.10x) 13,194 (1.02x) 685 (1.08x) 21,736 (1.10x) 1,360 (1.10x) 20,700 (1.11x) 1,476 (1.10x) 13,194 (1.00x) 3,257 (1.21x) 18,046 (1.05x) 7,251 (1.12x) 30,578 (1.04x)
TFB (Ours) 4 1,203 (3.56x) 10,372 (0.80x) 1,257 (1.99x) 11,966 (0.61x) 1,246 (1.01x) 11,202 (0.60x) 1,237 (0.92x) 10,344 (0.79x) 1,238 (0.46x) 10,376 (0.60x) 1,452 (0.22x) 16,340 (0.55x)
TFB (Ours) 8 628 (1.86x) 10,666 (0.83x) 731 (1.16x) 15,286 (0.77x) 702 (0.57x) 12,598 (0.68x) 634 (0.47x) 10,662 (0.81x) 642 (0.24x) 12,116 (0.70x) 1,015 (0.16x) 22,146 (0.75x)
TFB (Ours) 12 446 (1.31x) 12,064 (0.93x) 599 (0.94x) 18,204 (0.92x) 540 (0.43x) 14,310 (0.76x) 441 (0.32x) 11,370 (0.86x) 487 (0.18x) 13,410 (0.77x) 908 (0.13x) 25,220 (0.85x)
As shown in the table, for a small dataset, e.g., WG-S with $\sim$600 training examples, TFB can have a higher cost of time, especially when TFB is performed under the same batch size as the two baselines: almost 3x slower than BLoB. However, for a large dataset, e.g., BoolQ with $\sim$10,000 training examples, TFB Bayesianization process is almost 5x faster than BLoB using only half of the GPU memory. As TFB does not require gradient estimation, which significantly reduces the GPU memory, TFB can be further accelerated by increasing the batch size. When the batch size is increased from 4 to 12, TFB has a lower time and lower peak GPU memory usage compared to BLoB for almost all the datasets. TFB even has significantly better efficiency than standard LoRA fine-tuning, thanks to its training-free property.
Note that our efficiency analysis for TFB encompasses the complete parameter search process, while the reported metrics for other methods only include their final successful runs. This means the baseline measurements exclude significant hidden costs, particularly the computational resources required for hyperparameter tuning to determine optimal fine-tuning configurations. Therefore this comparison inherently favors the baseline method BLoB. Nevertheless, the fact that TFB still demonstrates superior efficiency in both time and memory consumption, even under these conditions that advantage the baselines, further underscores the benefits of its “training-free” approach. The results above also implies TFB’s flexibility, as it can adapt to limited-resource environments by trading time for reduced memory usage or vice versa, depending on operational requirements.
F.3 Anchor Dataset Size Analysis
We present the performance of BLoB-Mean + TFB on the ARC-Easy dataset with varying anchor dataset sizes ranging from 100 to 2000 in Figure 3. Initially, we hypothesized a negative correlation between the performance variance and the anchor dataset size. However, as shown in the figure, across experiments with three different random seeds, neither the performance variance nor the average performance exhibits a significant correlation with the anchor dataset size. This suggests that TFB is robust to the size of the anchor dataset.
Figure 3: Performance of BLoB-Mean + TFB with different size of anchor dataset on the ARC-Easy dataset. The shaded area represents the standard deviation of results across three random seeds, indicating that TFB is not sensitive to the size of the anchor dataset.
F.4 TFB Beyond Training Data as Anchor Dataset
In this section, we designate a portion of the unlabeled testing dataset as the anchor dataset, simulating a scenario where partial user input is accessible. The anchor dataset size remains fixed at 500 across all datasets. As illustrated in Table 10, the performance variation across different data sources for the anchor dataset is minimal, indicating that the choice of data source for the anchor dataset has negligible impact.
F.5 TFB Beyond the NLL Metric
We report the additional results of TFB when using Accuracy (ACC) as the evaluation metrics $l$ in Table 9. In our implementation, we adopt the change ratio of classification results as an accuracy-based evaluation metric ($l_{\text{ACC}}$) for the unsupervised anchor dataset. Comparing the two evaluation metrics ($l_{\text{ACC}}$ vs $l_{\text{NLL}}$) in Table 9, we observe comparable performance across all datasets. In some cases, accuracy-based evaluation ($l_{\text{ACC}}$) even yields slightly better results. For instance, BLoB-Mean+TFB achieves lower ECE on several datasets when using $l_{\text{ACC}}$. However, we adopt NLL as the primary evaluation metric in Table 1 since it better aligns with our theoretical framework in Theorem 4.2.
F.6 TFB Beyond the Low-Rank Isotropic Gaussians
In Sec. 5.3, we compare TFB with two alternative Gaussian distribution families that are controlled by a single parameter $\sigma_{q}$:
  • Constant Standard Deviation Matrix (C-STD): given $\sigma_{q}$, the C-STD’s variational distribution of the weight matrix $q(\operatorname{vec}({\bm{W}}))={\mathcal{N}}(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{q},{\bm{\Sigma}}_{q})$ where ${\bm{\mu}}_{q}={\bm{W}}_{0}+{\bm{B}}{\bm{A}}$ (same as TFB) and ${\bm{\Sigma}}_{q}=\sigma_{q}^{2}{\bm{I}}_{n}\otimes[{\bm{B}}{\bm{B}}^{\top}]$.
C-STD’s covariance matrix ${\bm{\Sigma}}_{q}$ is derived through Lemma D.1:
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ (37)
$=[{\bm{I}}_{n}\otimes{\bm{B}}][\sigma_{q}^{2}{\bm{I}}_{rn}][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ (38)
$=\sigma_{q}^{2}[{\bm{I}}_{n}\otimes{\bm{B}}][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ (39)
$=\sigma_{q}^{2}{\bm{I}}_{n}\otimes[{\bm{B}}{\bm{B}}^{\top}].$ (40)
We report the additional results comparing TFB with other approximate families Gaussians (FR and C-STD as discussed in Sec. 5.3) when using Accuracy as the evaluation metrics $l$ in Table 11. When the evaluation metric is set to Accuracy, the advantage of TFB becomes more significant compared to the results shown in Table 2. TFB with low-rank isotropic Gaussian as the variational distribution demonstrates superior calibration performance compared to both FR and C-STD variants while maintaining competitive accuracy. For ECE, TFB achieves better results across most datasets, with notable improvements on in-distribution tasks: 8.78% on WG-S (vs. 12.06% for FR and 11.61% for C-STD) and 1.28% on BoolQ (vs. 3.26% for FR and 2.65% for C-STD). Similarly for NLL, TFB consistently outperforms or matches the baseline variants, particularly on WG-S (0.55 vs. 0.63 for FR and 0.61 for C-STD) while preserving comparable accuracy scores. These results suggest that TFB’s approach to variance modeling is more effective than both full-rank isotropic and constant standard deviation alternatives.
F.7 TFB Beyond the Llama3.1-8B Backbone
The results demonstrate that TFB consistently improves model calibration across different backbones while maintaining competitive accuracy. Specifically, for Llama2-7B, TFB reduces the ECE from 4.50% to 1.24% on the combined dataset while preserving the accuracy (81.32% vs 81.41%). Similar improvements are observed with Llama3-8B, Llama3.1-8B, and Mistral-7B-v0.3, where TFB achieves better calibration than both the full training and early stopping baselines without sacrificing performance, suggesting its effectiveness as a general approach for enhancing LLM calibration.
F.8 TFB Beyond the Naive LoRA
  • VeRA [35]: Uses shared low-rank matrices ${\bm{B}}$ and ${\bm{A}}$ across layers, with layer-specific trainable scalar vector ${\bm{d}}$ and bias vector ${\bm{b}}$. Concretely, the parameterization of VeRA’s updated weight matrix ${\bm{W}}$ is modeled as:
    Hence after the fine-tuning of VeRA, we can easily regroup the weight matrices into $\{{\bm{B}}^{\prime}=[\operatorname{diag}({\bm{b}})]{\bm{B}}[\operatorname{diag}({\bm{d}})],{\bm{A}}^{\prime}={\bm{A}}\}$, and apply the TFB Bayesianization scheme illustrated in Algorithm A.
  • PiSSA [41]: Employs an alternative initialization scheme while maintaining LoRA’s parameterization and training procedure. Hence the TFB process for PiSSA is trivial.
The results in Table 13 show that TFB consistently improves calibration across different LoRA variants while preserving model performance. Notably, when applied to the standard LoRA, TFB significantly reduces the ECE from 4.74% to 1.05% on the combined dataset with minimal impact on accuracy (86.45% vs 86.70%). Similar improvements are observed with VeRA and PiSSA variants, where TFB achieves better calibration (reducing ECE to 1.44% and 1.17% respectively) while maintaining comparable accuracy levels. These results demonstrate that TFB can effectively enhance model calibration across different LoRA architectures without compromising their performance.
F.9 Improving the Inference-Time Efficiency of TFB
We report the detailed performance of last-layer TFB (LL TFB) in Table 14. As indicated in the table, with only $N=10$ samples, last-layer Bayesianization provides a less effective uncertainty estimation compared to full-model Bayesianization. However, increasing the number of samples to $N=100$ significantly enhances the posterior estimation, allowing last-layer Bayesianization to achieve better accuracy. This improvement further allows it to outperform the full-model Bayesianization in terms of NLL across most datasets.
F.10 Additional Results on Llama2-7B
We report the detailed performance of TFB applied to the Llama2-7B pre-trained weights in Table 15. The performance change tolerance $\epsilon$ is set adaptively to either 1% or 0.5%, depending on the checkpoint’s overfitting characteristics. To determine the optimal $\sigma_{q}^{*}$, we conduct parallel experiments with eight values of $\sigma_{q}\in$ [0.01, 0.015, 0.02, 0.025, 0.03, 0.035, 0.04, 0.05] using a single random seed. We construct an approximate function $\widehat{\sigma}_{q}(p)$ through piecewise linear interpolation of the observed performance and estimate $\sigma_{q}^{*}\approx\widehat{\sigma}_{q}(p_{0}-\epsilon)$. Similar to other baseline methods, the final results of TFB are reported as averages across three random seeds using $\sigma_{q}^{*}$.
In-Distribution (IND) Results. We observed several key patterns from the IND Datasets results. For example, the MLE baseline shows relatively strong accuracy but suffers from high ECE values (e.g., 29.83% on WG-S), indicating significant overconfidence. This aligns with the common challenge of LLM overconfidence during conventional fine-tuning.
TFB applied to BLoB-Mean demonstrates strong overall performance across the IND datasets, achieving the highest accuracy on several datasets (69.94% on WG-S, 70.72% on ARC-C, and 86.74% on ARC-E). More importantly, it achieves this while maintaining lower ECE values compared to methods like MCD and ENS, suggesting better calibrated predictions. The method also shows strong NLL performance, with values consistently among the lowest across datasets (0.62 for WG-S, 0.86 for ARC-C).
In summary, TFB consistently enhances the performance of baseline methods (MLE, MAP, and BLoB-Mean) across different evaluation scenarios, with notable improvements in both accuracy and calibration metrics. The improvements are particularly evident in the significant ECE reductions (e.g., from 29.83% to 16.26% for MLE on WG-S) while maintaining or improving accuracy, with the most substantial gains observed when TFB is combined with BLoB-Mean, achieving both the highest accuracy and lowest ECE values across most datasets.
Out-of-Distribution (OOD) Results. The OOD evaluation reveals interesting patterns across both smaller and larger distribution shifts. For smaller shifts (ARC-C and ARC-E), BLoB-Mean with TFB maintains strong performance, achieving 70.38% and 80.16% accuracy respectively, while keeping ECE values low (12.28% and 8.07%). This suggests robust generalization under moderate distribution shifts.
For larger shifts (Chem and Phy datasets), we see a more significant performance degradation across all methods, as expected. However, BLoB-Mean with TFB still maintains competitive performance, achieving 42.67% accuracy on Chem and 30.67% on Phy, while maintaining reasonable calibration metrics. The method’s NLL values (1.35 and 1.46 respectively) remain competitive with other approaches, indicating relatively well-calibrated uncertainty estimates even under substantial distribution shifts.
Notable is the consistently strong performance of the BLoB variants (both w/ and w/o TFB) across different metrics and datasets, suggesting that this approach offers a robust framework for both in-distribution and out-of-distribution scenarios. The results demonstrate that the method successfully balances the trade-off between accuracy and calibration, particularly evident in the out-of-distribution scenarios where maintaining both aspects becomes more challenging.